OpenStudy (anonymous):

Question 6 (Multiple Choice Worth 3 points) Solve: x + 10 = −x + 12 8 −4 4 Question 7 (Multiple Choice Worth 3 points) Solve: 2(x + 6) − (x + 1) = 16 3 5 2 1 Question 8 (Multiple Choice Worth 3 points) Solve: 2 − x = 13 −11 11 −15 15 Question 9 (Multiple Choice Worth 3 points) Solve: 4(x + 1) = −12 −2 4 −3.25 −4

OpenStudy (unklerhaukus):

\[x + 10 = −x + 12\] take away ten form both sides , and add x to both sides

OpenStudy (unklerhaukus):

after you have done that divide both sides by two

OpenStudy (unklerhaukus):

what'd ya get?

OpenStudy (anonymous):

6

OpenStudy (unklerhaukus):

you should have an equation

OpenStudy (anonymous):

2x=2??

OpenStudy (anonymous):

sooooooooooooooooooooooooooooo???????????????????

OpenStudy (unklerhaukus):

good. now divide both sides by two

OpenStudy (anonymous):

and i get 1

OpenStudy (unklerhaukus):

is that the third option for Question 6?

OpenStudy (anonymous):

no its 1/2

OpenStudy (anonymous):

Light is the manifestation of a "microscopic" phenomenon.

OpenStudy (unklerhaukus):

well we can check our solution \[x + 10 = −x + 12\]\[x=1\] \[1 + 10 = −1 + 12\] \[\qquad11=11\]\(\text {true}\therefore x=1, \) is a solution to \(x + 10 = −x + 12\)

OpenStudy (unklerhaukus):

Question 7\[ 2(x + 6) − (x + 1) = 16\] first expand the brackets on the left hand side

OpenStudy (anonymous):

2x+6-(x+1=16

OpenStudy (unklerhaukus):

try again

OpenStudy (anonymous):

2x+12 @UnkleRhaukus

OpenStudy (unklerhaukus):

that is the proper expansion of the first brackets keep going